Mathematics 133 – Intensive Calculus for Science 1 Quiz 5 – October 14, 2005

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. Calculators allowed on this quiz.

A) If f(x) = 2x^{2} − x + 5, use the limit definition to compute the derivative function f′(x).

B) The first of the following graphs shows y = f(x). Use it to make a “qualitative” sketch of y = f′(x) on the second set of coordinate axes, showing all points where f′(x) = 0, where f′(x) > 0 and where f′(x) < 0.